# Training Machine Learning models at the Edge: A Survey

### v2

@@ V-B3 Meta-learning
Meta learning focus on enhancing a model’s ability to learn new tasks quickly and effectively. Unlike traditional learning paradigms that optimize for a specific task, meta learning trains models to learn from a diverse set of tasks, thereby enabling them to generalize knowledge and adapt rapidly to novel tasks with minimal data [141]. By exposing models to various learning scenarios, meta learning equips them with transferable skills, such as recognizing patterns and adapting to new contexts.
@@
In the context of edge learning, the application of meta learning introduces a transformative approach to address the challenges posed by limited data availability and resource constraints [142, 143]. [144] proposes adaptation-aware network pruning, a model pruning method designed to work with existing meta learning methods to achieve fast adaptation on edge devices, while [136] proposes a continual meta-learning approach with bayesian graph neural networks that mathematically formulates meta-learning as continual learning of a sequence of tasks, and p-Meta was introduced in [142], and aims to achieve faster generalization to unseen tasks and enforces structure-wise partial parameter updates to support memory-efficient adaptation. Meta learning can also be used in accordance to other techniques such as federated learning [135, 145], or [95] that integrates reinforcement learning models trained by multiple edge devices into a general model based on a meta-learning approach, in order to create FedMC, a generalized federated reinforcement learning framework based on a meta-learning approach.
V-C Model Compression based Techniques
@@
Within this section, we embark on an exploration of model compression techniques aimed to streamline the training of ML models at the edge. As traditional deployment and inference solutions have embraced knowledge distillation, quantization, and model pruning to streamline model execution on resource-constrained devices, a notable shift is observed towards employing these techniques for reducing the complexity of ML models for the training in the edge, making these techniques helpful for the training phase as well.
V-C1 Knowledge distillation
@@
Knowledge Distillation (KD) in deep learning is a process whereby a small or student neural network is trained to emulate the knowledge and predictive capabilities of a larger or teacher network. This technique serves as a means to transfer the expertise and generalization capabilities of a complex model to a simpler one, thereby enhancing inference efficiency and reducing computational demands [146]. The underlying principle involves the student network learning not only from ground truth labels but also from the soft, probabilistic outputs of the teacher network, thereby capturing finer details and nuances in the data [146]. In the context of edge learning, knowledge distillation is usually used to reduce the size and complexity of a large neural network, to simplify its training in limited resources devices. Therefore, knowledge distillation is well suited to be used in collaboration with other techniques such as federated learning [25, 98], split learning [147] or incremental learning [137]. However, distillation is also used independently of other techniques [148, 149].
@@
In the context of using knowledge distillation with FL on edge devices, recent trends indicate great potential in combining the two techniques [25], and considerable approaches have been developed in that regard, from an Attack-Resistant Federated Edge Learning method with dataset distillation [150] to speech recognition tasks [92] or keyword spotting [151]. Mix2FLD [152] is another method that combine KD and FL, while [99] use both KD and Quantization to train Federated Learning models on Edge Devices. Other hybrid methods involving FL and KD have been proposed [153, 154], some of them are explored in details in Wu and Al’s Survey of Knowledge Distillation in Federated Edge Learning [25]. Other Distributed learning methods can benefit from KD as well, for instance [147] introduces a spatio-temporal distillation method for split learning for a tiny server in order to alleviate the frequent communication costs that happen when communicating from the server to edge devices, and [155] introduces a distributed distillation algorithm where devices communicate and learn from soft-decision outputs, which are inherently architecture-agnostic and scale only with the number of classes in order to alleviate the communication costs from transmitting model weights in the network and improve the inclusion of devices with different model architectures. Finally, KD has been used as a standalone technique in an edge learning context, for recommendation systems [156, 157, 148], Edge Cardiac Disease Detection [158] and on-device Deep Reinforcement Learning [159]. KD was, additionally, used with multiple variants, including dataset distillation techniques [160, 161] and knowledge transfer [162, 149].
V-C2 Quantization
@@
Quantization in deep learning refers to the process of reducing the precision of numerical values representing model parameters or activations, typically from floating-point to fixed-point or integer representations, in order to balance the act of maintaining an acceptable level of model accuracy while significantly reducing the memory and computational requirements [163]. This computational optimization technique is pivotal in mitigating the resource-intensive demands of deep neural networks, rendering them more amenable for resource-constrained hardware platforms, such as edge devices and embedded systems[163]. There are two types of quantization: Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ). In QAT, the quantized model is fine-tuned using training data in order to adjust parameters and recover accuracy degradation or perturbation introduced by the quantization, while in the less expensive PTQ the pretrained model is calibrated using calibration data before the quantization which is based on the calibration result [163].
@@
Although quantization techniques are usually used to optimize machine learning models before deployment for edge inference [164, 121], it can be used in edge learning in order to simplify the fine-tuning of large models in the edge [31, 165]. Similarly to KD, quantization is often alongside other techniques such as federated learning [166], transfer Learning [31], incremental learning [167] or with other types of techniques [168]. Quantization is used with FL particularly extensively, from one-bit quantization [169, 170, 171] to hierarchical FL [100] and other FL based approaches such as [172, 173, 174, 175, 176, 99]. Other Quantization based methods for training ML models in the edge include, Quantization-Aware Scaling (QAS), a method proposed in [31], to automatically scale the gradient of tensors with different bit-precisions without requiring any fine-tuning, the method was used alongside the Tiny Training Engine and Sparse Update introduced in the same paper [31]. In [177], investigations were made on on-device few shot learning for audio classification, and it was found that quantization helps further in reducing the resource requirements for the training. In the Holmes optimizer [178], quantization is used in an ingenious way to improve the accuracy by effectively combining different types of quantization (a limitation on the number of quantization bits, fixed-point number, and logarithmic quantization) techniques.
V-C3 Model Pruning
@@
Model pruning is a technique used to reduce the size of ML models by removing certain parts of the model, such as model parameters, nodes in a decision tree [179] or weight matrices in transformer-based models [180]. Similarly to quantization, model pruning is commonly used in the edge to improve the efficiency of inference and reduce computational resources required for the deployment of ML models [181]. However, much like quantization, model pruning has also shown great potential in edge learning, by reducing the size of the ML model before fine-tuning it on edge devices, particularly used in conjunction with other techniques such as federated learning [174, 182, 183, 184], incremental learning [131, 139] or meta-learning [144].
@@
Similar to knowledge distillation and quantization, federated learning emerges as the most prominent technique when combined with model pruning for EL. Noteworthy is PruneFL [183], an approach aimed at minimizing communication and computation overhead while reducing training time through adaptive model size adjustment during FL. PruneFL employs model pruning, starting with an initial pruning stage at a selected client, followed by subsequent pruning iterations during FL. Additionally, [184] introduce model pruning for wireless Federated Learning to scale down neural networks, while [182] employ an adaptive dynamic pruning approach to prevent overfitting by slimming the model through the dropout of unimportant parameters. Other approaches that use model pruning on the edge include [185] that uses model pruning in the context of On-Device Personalization for an activity recognition system, and Deeprec [186] which leverages model pruning and embedding sparsity techniques to reduce computation and network overhead. OmniDRL [187], is a Deep Reinforcement Learning based approach on edge devices, incorporates weight pruning in each learning iteration to achieve a high weight compression ratio. Finally, [188] explores the reduction in memory footprint for further pruning during the training phase of BitTrain, a bitmap memory efficient compression technique for training on edge devices.
V-D Optimization and Acceleration based Techniques
@@
In this section, we explore some of the multiple other techniques that do not necessarily fall into a specific category and are used to optimize or provide more optimized alternatives to machine learning models’ computation efficiency, memory footprint, energy consumption, etc, which makes them more suited for edge learning.
V-D1 Binary neural networks
@@
Binary Neural Networks (BNNs) are Deep Neural Networkss (DNNs) that use binary values (-1 or 1) instead of floating-point numbers for weights and activations. BNNs are attractive for resource-constrained devices due to their ability to compress DNNs [189]. Similar to other techniques such as quantization and model pruning, BNNs are considered to be good candidates for edge inference due to their extreme compute and memory savings over higher-precision alternatives [190]. However, multiple approaches were proposed to leverage BNNs compute and memory efficiency for edge learning, such as [191] that proposes a hybrid quantization of a Continual Learning model with a BNN as its backbone, or [192] that propose a BNN model based on an MRAM array with Ternary Gradients for both training and inference on the edge, other BNN based approaches for edge learning include [193, 194, 190].
V-D2 Spiking neural networks
@@
Spiking neural networks (SNNs) are another type of DNNs that is promising for the edge, SNNs communicate between neurons using events called spikes [195] and are known for their asynchronous and sparse computations, which result in decreased energy consumption [196, 197], and this make them well suited for edge computing especially for energy limited devices [198]. Training ML models at the edge using SNNs had gathered some traction, [199] propose FL-SNN, a cooperative training through FL for networked on-device SNNs, [200] present a memristor spiking neuron and synaptic trace circuits for efficient on device learning, and [143] integrate a meta-learning approaches with SNNs that enables lifelong learning on a stream of tasks, and relies on local backpropagation-free nested updates. SNNs can also be trained at the edge using event-driven, power and memory-efficient local learning rules, such as spike-timing-dependent plasticity (STDP), such as [201] that use STDP and Synaptic Normalization for unsupervised on-device learning. There are other approaches that leverage SNNs for EL, including [202, 203, 204].
V-D3 Forward-Forward Algorithm
@@
The backpropagation algorithm is essential for training neural networks, but recent studies have proposed alternatives to the algorithm when the available resources are limited, such as the Forward-Forward (FF) [205] algorithm that aims to replace the forward and backward passes of backpropagation by two forward passes that operate in the same way as each other on different data and with opposite objectives. A positive pass operates on real data and adjusts the weights to increase the goodness in every hidden layer, while a negative pass operates on ”negative data” and adjusts the weights to decrease the goodness in every hidden layer [205]. Some other studies have tried to adapt the Forward-Forward algorithm to edge devices, such as µ-FF [206] that propose an on-device variation of the original FF on microcontroller units which tackles the training process with a multivariate Ridge regression approach and allows finding closed-form solution by using the Mean Squared Error (MSE) as loss function. And finally, [207] investigates the improvements in terms of complexity and memory usage brought by PEPITA [208] and the Forward-Forward Algorithm, the results show that the Forward-Forward algorithm reduces memory consumption by 40% on average, however it involves additional computation at inference that, can be costly on microcontrollers.
V-D4 Other techniques
@@
In this section, we will explore some other techniques used to optimize ML models for training in the edge. One such techniques is Data Booleanization used in [209] which proposes a novel approach towards low-energy booleanization. MiniLearn [210] on the other hand, enables re-training of DNNs on resource-constrained IoT devices and allows IoT devices to re-train and optimize pre-trained, quantized neural networks using IoT data collected during the deployment of an IoT device. While [211], propose the use of Echo State Networks for anomaly detection on-the-edge in aerospace applications. Tiny Training Engine (TTE) [31] is a lightweight training system, introduced alongside Sparse Update, a technique that skip the gradient computation of less important layers and sub-tensors, and a Quantization-Aware Scaling to stabilize 8-bit quantized training, TTE enables on-device training of convolutional neural networks under 256KB of SRAM and 1MB Flash without auxiliary memory [31].
@@
[212] introduces a novel reduced precision optimization technique for On-Device Learning primitives on MCU-class devices with a specialized shape transform operators and matrix multiplication kernels, accelerated with parallelization and loop unrolling for the backpropagation algorithm. POET [213], allow for the training of large neural networks on memory scarce and battery-operated edge devices, with Integrated Rematerialization and Paging POET reduce the memory consumption of backpropagation and could fine-tune both ResNet-18 and BERT within edge devices memory constraints. Finally [168] proposes a novel rank-adaptive tensor-based tensorized neural network mode for on-device trainging with ultra-low memory usage.
V-E Comparison of techniques used in edge learning
@@
In this section we will compare the different families of techniques used to train ML models in the edge that we explored in the previous part of this paper (sections V-A, V-B, V-C, V-D), we will compare the different families based on two factors : 1. The number of academic contributions of each different technique family and their evolution over the years; 2. The techniques’ potential of answering the different needs and requirements particular with edge learning.
V-E1 Comparison of the usage of the different techniques
@@
We will first start by making a comparison of the different EL techniques’ over the years, by analysing their academic contributions. The figure 4 presents the number of papers for each technique, while the figure 5 references the number of papers (in a logarithmic scale) per technique per year. In order to get the necessary data, we used the advanced search feature of Scopus  [^1] Scopus : https://www.scopus.com/ where we searched the terms ”edge learning”, ”training/learning on the edge/mobile devices”, ”on-device training/learning” and ”on-device adaptation” as well as the relevant keywords for each technique (”Federated Learning”, ”Split Learning”, etc), in the Title, Keywords and Abstract, we made sure to exclude the surveys, conference reviews, books and notes as we are only interested in approaches to train ML models in the edge using the aforementioned techniques. For each family of techniques, we manually reviewed and removed papers that either didn’t provide a contribution, or weren’t about training ML models on edge devices with said technique, despite checking the keywords. Finally, we added a few papers we found manually and either were not available in Scopus or didn’t check the keywords but were relevant in the context of the analysis. The final number of papers associated with the analysis was 511 papers, and some of these papers were briefly covered in V. Note that multiple techniques can be used in a single paper, therefore the total count of techniques in the figures 4 and 5 will exceed 511.
Figure 4: Overall trend of techniques used to train ML models in the edge
@@
The analysis of Figure 4 reveals that Federated Learning (FL) is the dominant approach for training ML models in edge environments, given the resource constraints of edge devices. This dominance is expected, as distributed learning methods that capitalize on the collective computing power of multiple devices are deemed more practical and efficient in edge settings. Moreover, we anticipate that this trend will persist and expand to include Split Learning (SL), another promising distributed learning technique. Other methods, including Incremental Learning (IL), Transfer Learning (TL), Model Compression Techniques (e.g., Quantization, Knowledge Distillation (KD)), although consistently employed, lag behind FL in terms of popularity. Furthermore, techniques such as Swarm or Gossip Learning have received scant attention. A closer examination of Figure 5 unveils a remarkable surge in the number of publications focused on edge learning over the past six years. Notably, there has been a steady rise in the adoption of FL and Split Learning from 2018 to 2023, aligning with our forecast of a burgeoning trend favoring these two techniques. In contrast, the use of techniques like IL and TL has been more steady during the same period, and meta-learning despite being employed consistently from 2019 to 2022 received little attention in 2023. Finally, for the model compression techniques while quantization and KD experienced a small rise in popularity over the years, model pruning has exhibited greater fluctuation during that period.
As previously discussed, various approaches have been proposed that integrate multiple techniques to mitigate the limitations of individual methods and capitalize on their respective strengths. Examples of such approaches include [94, 135, 136, 122]. To provide a clearer illustration of the relationships between these various techniques, a heatmap depicting the intersection of their usage is presented in Figure 6. This visual representation allows for a more comprehensive understanding of the synergies and overlap between different approaches. We can note that FL has been combined the most with other techniques, which is expected considering the overwhelming number of FL contributions to the edge (see Figure 4). Additionally, Model Compression techniques such as KD, model pruning and quantization are also used together with other techniques as discussed in Sections (V-C1 V-C2 V-C3). Finally, we can note other collaborations between the different families, and we expect this trend to continue in the future.
@@
V-E2 Comparison based on the requirements and needs of EL
As discussed in Section IV, several requirements must be met for edge learning, which can function as imprecise measures for assessing the viability of the families of techniques covered in earlier sections. After reevaluating the prerequisites, we eliminated ”Labelled Data Independence” from the comparison, because it is more relevant to the type of machine learning employed and the availability of an autolabeling process than to the approaches under consideration. Additionally, we do not consider ”Task-specific metrics and performance” since that covers multiple metrics used in specific tasks; nevertheless, we introduce a ”High Performance” measure to estimate roughly if the strategies positively or negatively affect the performances. For instance, model compression techniques often lower model performance, whereas incremental and meta-learning typically boost it. Consequently, we employ six distinct measures: ”Computational Efficiency,” ”Memory Footprint,” ”Low Energy Consumption,” ”Quick Training Time,” ”Optimized Bandwidth,” and ”High Performance.” Table II compares the various families of techniques against these requirements.
The outcomes, as depicted in Table II, should be interpreted as informal and approximate assessments aimed at providing a broad understanding of the general strengths and weaknesses of the compared techniques within the context of edge learning (EL). A checkmark (”✓”) implies that, in general, the technique offers assistance or advantages when applied to EL with respect to the specified requirement. Conversely, a cross mark (”✗”) indicates that, in general, the requirement represents a weakness of the technique in the EL context. It is important to note that techniques like Federated Learning (FL), quantization, and others have various variants and specific approaches that influence how these methods align with the requirements. For instance, while quantization techniques may result in a minor decrease in performance in most scenarios [214], leading to an ”✗” check in the ”High Performance” column. Certain specific approaches to quantization may exceptionally yield no performance loss or even improvements, as demonstrated in studies such as [178]. Therefore, it is important to note that the assessment provided by these symbols should be considered as rough estimates, as the effectiveness of a technique, and how it fares against a specific requirement, can vary depending on diverse factors such as variant versions, implementation details, use cases, tasks, and hardware platforms. Accordingly, the table II offers a general overview rather than a definitive judgment on the suitability of each technique for every situation.

### v3

@@ III-B3 Meta-learning
Meta learning focus on enhancing a model’s ability to learn new tasks quickly and effectively. Unlike traditional learning paradigms that optimize for a specific task, meta learning trains models to learn from a diverse set of tasks, thereby enabling them to generalize knowledge and adapt rapidly to novel tasks with minimal data [151]. By exposing models to various learning scenarios, meta learning equips them with transferable skills, such as recognizing patterns and adapting to new contexts. For further insight into meta learning, we recommend consulting [152, 153].
@@
In the context of edge learning, the application of meta learning introduces a transformative approach to address the challenges posed by limited data availability and resource constraints [154, 155][156] proposes adaptation-aware network pruning, a model pruning method designed to work with existing meta learning methods to achieve fast adaptation on edge devices, while [146] proposes a continual meta-learning approach with bayesian graph neural networks that mathematically formulates meta-learning as continual learning of a sequence of tasks, and p-Meta was introduced in [154], and aims to achieve faster generalization to unseen tasks and enforces structure-wise partial parameter updates to support memory-efficient adaptation. Meta learning can also be used in combination with other techniques such as FL [145, 157], or [103] which integrates reinforcement learning models trained by multiple edge devices into a general model based on a meta-learning approach, in order to create FedMC, a generalized federated reinforcement learning framework based on a meta-learning approach.
III-C Model Compression based Techniques
@@
This section explores model compression techniques, which aim to streamline the training of ML models at the edge. As traditional deployment and inference solutions have embraced knowledge distillation, quantization, and model pruning to accelerate model execution on resource-constrained devices, a notable shift is observed towards employing these techniques for reducing the complexity of ML models for the training in the edge, making these techniques helpful for the training phase as well.
III-C1 Knowledge distillation
@@
Knowledge distillation in deep learning is a process whereby a small or student neural network is trained to emulate the knowledge and predictive capabilities of a larger or teacher network. This technique serves as a means to transfer the expertise and generalization capabilities of a complex model to a simpler one. As a result, inference efficiency is enhanced, and computational demands are reduced. The underlying principle involves the student network learning not only from ground truth labels but also from the soft, probabilistic outputs of the teacher network, thereby capturing finer details and nuances in the data [158]. For a deeper dive into knowledge distillation, readers can refer to the following survey  [158]. In the context of edge learning, knowledge distillation is usually used to reduce the size and complexity of a large neural network, to simplify its training in limited resources devices. Therefore, knowledge distillation is well suited to be used in collaboration with other techniques such as federated learning [28, 106], split learning [159] or incremental learning [147]. However, distillation is also used independently of other techniques [160, 161].
@@
The integration of knowledge distillation with FL on edge devices has shown promising results, with recent trends indicating great potential in combining the two techniques [28]. Several approaches have been proposed, including attack-resistant FL methods [162], speech recognition tasks [96] or keyword spotting [163]. Mix2FLD [164] is another method that combines knowledge distillation and FL. Meanwhile, [107] use both distillation and quantization to train FL models on edge devices. Several other hybrid approaches combining FL and knowledge distillation have been explored [165, 166]. These approaches are further discussed in Wu et al.’s survey on knowledge distillation in federated edge learning [28]. Knowledge distillation can also be applied to other distributed learning methods, such as [159], which introduces a spatio-temporal distillation method for split learning for a tiny server in order to alleviate the frequent communication costs that happen when communicating from the server to edge devices. [167] introduces a distributed distillation algorithm where devices communicate and learn from soft-decision outputs, which are inherently architecture-agnostic and scale only with the number of classes in order to alleviate the communication costs from transmitting model weights in the network and improve the inclusion of devices with different model architectures. Finally, knowledge distillation has been used as a standalone technique in an edge learning context, for recommendation systems [168, 169, 160], edge cardiac disease detection [170] and on-device deep reinforcement learning [171]. Knowledge distillation was, additionally, used with multiple variants, including dataset distillation techniques [172, 173] and knowledge transfer [174, 161].
III-C2 Quantization
@@
Quantization in deep learning refers to the process of reducing the precision of numerical values representing model parameters or activations, typically from floating-point to fixed-point or integer representations, in order to balance the act of maintaining an acceptable level of model accuracy while significantly reducing the memory and computational requirements [175]. This computational optimization technique is pivotal in mitigating the resource-intensive demands of deep neural networks, rendering them more amenable for resource-constrained hardware platforms, such as edge devices and embedded systems. There are two types of quantization: quantization-aware training and post-training quantization. In quantization-aware training, the quantized model is fine-tuned using training data in order to adjust parameters and recover accuracy degradation or perturbation introduced by the quantization. In contrast, post-training quantization is a less expensive approach, where the pretrained model is quantized and its weight adjusted without any fine tuning [175]. To gain a more complete understanding of quantization techniques, readers are advised to consult [175]
@@
Quantization techniques are not only used to optimize machine learning models before deployment for edge inference [176, 131], but also in edge learning to simplify the fine-tuning of large models [44, 177]. Similarly to knowledge distillation, quantization is often used alongside other techniques such as FL [178], transfer Learning [44], incremental learning [179] or with other types of techniques [180]. Quantization is used with FL particularly extensively, including one-bit quantization [181, 182, 183] and hierarchical FL [108]. Other FL-based approaches that utilize quantization include [184, 185, 186, 187, 188, 107]. Other quantization-based methods for training ML models in the edge include quantization-aware scaling, which was proposed in [44]. This method automatically scales the gradient of tensors with different bit-precisions without requiring any fine-tuning, and was used alongside a tiny training engine and sparse updates. Investigations in [189] showed that quantization helps further in reducing the resource requirements for the training on on-device few shot learning for audio classification. The Holmes optimizer [190] uses quantization to improve the accuracy by combining different quantization techniques, such as limiting quantization bits, fixed-point numbers, and logarithmic quantization.
III-C3 Model Pruning
@@
Model pruning is a technique used to reduce the size of ML models by removing certain parts of the model, such as model parameters, nodes in a decision tree [191] or weight matrices in transformer-based models [192]. Similarly to quantization, model pruning is commonly used in the edge to reduce the computational resources required for the inference of ML models [193]. Model pruning has also shown great potential in edge learning, where it reduces the size of ML model before fine-tuning on edge devices. This technique is particularly effective when used in conjunction with other methods, such as FL [186, 194, 195, 196], incremental learning [141, 149] or meta-learning [156].
@@
Similar to knowledge distillation and quantization, FL emerges as the most prominent technique when combined with model pruning for edge learning. Noteworthy is PruneFL [195], an approach aimed at minimizing communication and computation overhead while reducing training time through adaptive model size adjustment during FL. PruneFL employs model pruning, starting with an initial pruning stage at a selected client, followed by subsequent pruning iterations during FL. Other approaches that combine model pruning with FL include [196] which introduces model pruning for wireless FL to scale down neural networks. Meanwhile, [194] employ an adaptive dynamic pruning approach to prevent overfitting by slimming the model through the dropout of unimportant parameters. In addition, several approaches use model pruning on edge devices. For instance, [197] uses model pruning in the context of on-device personalization for an activity recognition system, and Deeprec [198] leverages model pruning and embedding sparsity techniques to reduce computation and network overhead. Furthermore, OmniDRL [199], a deep reinforcement learning based approach on edge devices, incorporates weight pruning in each learning iteration to achieve a high weight compression ratio. Finally, [200] explores the reduction in memory footprint for further pruning during the training phase of BitTrain, a bitmap memory efficient compression technique for training on edge devices.
III-D Optimization and Acceleration based Techniques
@@
In this section, we explore some other techniques that don’t fit neatly into a specific category and are used to optimize or provide more optimized alternatives to machine learning models’ enabling them to be more suitable for edge learning.
III-D1 Binary neural networks
@@
Binary Neural Networks (BNNs) are deep neural networks that use binary values (-1 or 1) instead of floating-point numbers for weights and activations. BNNs are attractive for resource-constrained devices because of their ability to compress deep neural networks [201]. BNNs share similarities with other techniques, such as quantization and model pruning, which are also considered good candidates for edge inference due to their extreme compute and memory savings over higher-precision alternatives [202]. However, BNNs’ compute and memory efficiency can also be leveraged for edge learning. For example, by proposing a hybrid quantization of a continual learning model [203], or by developing a model based on an MRAM array with ternary gradients for both training and inference on the edge [204]. Other BNN based approaches for edge learning include [205, 206, 202].
III-D2 Spiking neural networks
@@
Spiking neural networks (SNNs) are another type of deep neural networks that are promising for the edge. SNNs communicate between neurons using events called spikes [207] and are known for their asynchronous and sparse computations. These properties result in decreased energy consumption [208, 209], which makes them well suited for energy limited devices [210]. Training ML models at the edge using SNNs has gained some attention in recent years. For example, [211] proposes FL-SNN, a cooperative training through FL for networked on-device SNNs, while [212] presents a memristor spiking neuron and synaptic trace circuits for efficient on device learning. Other approaches include integrating meta-learning with SNNs for lifelong learning on a stream of tasks with local backpropagation-free nested updates [155], and using event-driven, power and memory-efficient local learning rules, such as spike-timing-dependent plasticity [213]. There are other approaches that leverage SNNs for edge learning, including [214, 215, 216].
III-D3 Forward-Forward Algorithm
@@
The backpropagation algorithm is essential for training neural networks, but recent studies have proposed alternatives to the algorithm when the available resources are limited. One such algorithm is the forward-forward [217] algorithm, which replaces the forward and backward passes of backpropagation by two forward passes that operate in the same way as each other on different data with opposite objectives. A positive pass operates on real data and adjusts the weights to increase the goodness in every hidden layer, while a negative pass operates on ”negative data” and adjusts the weights to decrease the goodness in every hidden layer [217]. To adapt the forward-forward algorithm to edge devices, researchers have proposed variations such as µ-FF [218], which uses a multivariate ridge regression approach and allows finding closed-form solution by using the mean squared error. Another study [219] investigates the improvements in terms of complexity and memory usage brought by PEPITA [220] and the forward-forward algorithm. the results show that the forward-forward algorithm reduces memory consumption by 40% on average, but involves additional computation at inference that, can be costly on microcontrollers.
III-D4 Other techniques
@@
In this section, we will explore some other techniques used to optimize ML models for training in the edge. One such technique is data booleanization, used in [221], which proposes a novel approach towards low-energy booleanization. MiniLearn [222] on the other hand, enables re-training of deep neural networks on resource-constrained IoT devices. This allows them to re-train and optimize pre-trained, quantized neural networks using IoT data collected during deployment. Another approach is the use of echo state networks for anomaly detection in aerospace applications [223]. Tiny training engine [44] is a lightweight training system, introduced alongside sparse update, a technique that skip the gradient computation of less important layers and sub-tensors, and a quantization-aware scaling to stabilize 8-bit quantized training. Tiny training engine enables on-device training of convolutional neural networks under 256KB of SRAM and 1MB flash without auxiliary memory [44].
@@
[224] introduces a novel reduced precision optimization technique for on-device learning primitives on MCU class devices with a specialized shape transform operators and matrix multiplication kernels, which is accelerated with parallelization and loop unrolling for the backpropagation algorithm. In addition, POET [225] allows for the training of large neural networks on memory scarce and battery-operated edge devices, with integrated rematerialization and paging. POET reduce the memory consumption of backpropagation, allowing the fine-tuning of both ResNet-18 and BERT within edge devices’ memory constraints. Finally, [180] proposes a novel rank-adaptive tensor-based tensorized neural network mode for on-device training with ultra-low memory usage.
III-E Comparison of techniques used in edge learning
@@
In this section we will compare the different families of techniques used to train ML models in the edge that we explored in the previous part of this paper (sections III-AIII-BIII-CIII-D), we will compare the different families based on two factors: (i) The number of academic contributions of each family and their evolution over the years; (ii) The techniques’ potential of answering the different needs and requirements particular to edge learning.
III-E1 Comparison of the usage of the different techniques
@@
We will first start by comparing the different edge learning techniques over the years by analyzing their academic contributions. Figure 3 shows the number of papers per technique per year on a logarithmic scale. We used the advanced search features of Scopus [^1] Scopus: https://www.scopus.com/ and Web of Science [^2] Web of Science: https://clarivate.com/products/scientific-and-academic-research/research-discovery-and-workflow-solutions/webofscience-platform/ to get the data. Using those search engines, we searched for the terms ”edge learning”, ”Edge Intelligence”, ”training/learning on the edge/mobile devices”, ”on-device training/learning” and ”on-device adaptation” as well as relevant keywords for each technique (”Federated Learning”, ”Split Learning”, etc.), in the title, keywords and abstract. We excluded surveys, books, and notes, as we are only interested in technical contributions to optimizing ML model training on the edge using the aforementioned techniques. Additionally, we manually reviewed each paper and removed papers that either didn’t provide a technical contribution, or weren’t about training ML models on the edge, despite containing relevant keywords. Finally, we added a few manually found papers that were not indexed in both Scopus and Web of Science or did not contain the relevant keywords, but were still relevant to our analysis. Note that we excluded the families of techniques that had less than 10 papers in total for edge learning. The excluded techniques are: swarm learning, gossip learning, forward-forward, BNNs and SNNs. The final number of papers associated with the analysis was 803 papers, and some of these papers were briefly covered in III. Note that multiple techniques can be used in a single paper (see Figure 4), therefore the total count of techniques in the Figure 3 will exceed 803. The cut-off date for the year 2024 is the 20th July 2024.
@@
The analysis of Figure 3 reveals that FL is the dominant approach for training ML models in edge environments, given the resource constraints of edge devices. This dominance is expected, as distributed learning methods that capitalize on the collective computing power of multiple devices are deemed more practical and efficient in edge settings. Moreover, we anticipate that this trend will persist and expand to include split learning, another promising distributed learning technique. Other methods, including incremental learning, transfer learning, Model Compression Techniques (e.g., quantization, knowledge distillation), although consistently employed, lag behind FL in terms of popularity. A closer examination of Figure 3 unveils a remarkable surge in the number of publications focused on edge learning over the past six years. Notably, there has been a steady rise in the adoption of FL and split learning, aligning with our forecast of a trend favoring these two techniques. In contrast, the use of techniques like incremental learning and transfer learning has been more steady during the same period, and meta-learning despite being employed consistently in previous years received less attention in 2023. Finally, for the model compression techniques while quantization and knowledge distillation experienced a small rise in popularity over the years, model pruning has exhibited greater fluctuation during that period.
As previously discussed, various approaches have been proposed that integrate multiple techniques to mitigate the limitations of individual methods and capitalize on their respective strengths. Examples of such approaches include [102, 145, 146, 132]. To provide a clearer illustration of the relationships between these various techniques, a heatmap depicting the intersection of their usage is presented in Figure 4. This visual representation allows for a more comprehensive understanding of the synergies and overlap between different approaches. The heatmap includes all the technique families discussed in Section III, except for swarm learning, gossip learning, and the forward-forward algorithm, which have not been combined with other techniques in the context of edge learning to our knowledge. We can note that FL has been combined the most with other techniques, which is expected considering the overwhelming number of FL contributions to the edge (see Figure 3). Furthermore, model compression techniques such as knowledge distillation, model pruning and quantization are also often used together with other techniques as discussed in Sections (III-C1 III-C2 III-C3). Finally, we can observe other collaborations between the different families, and we expect this trend to continue in the future.
@@
III-E2 Comparison based on the requirements and needs of edge learning
As outlined in Section II-C, there are several requirements that must be met for edge learning, which can function as imprecise measures for assessing the viability of the families of techniques covered in earlier sections for the training in the edge. Upon reevaluating the requirements, we excluded “labelled data independence” from the comparison, as it is more related to the type of ML employed or the availability of an autolabeling process rather than to the techniques being evaluated. Furthermore, we do not consider “task-specific metrics and performance” as this encompasses multiple metrics that vary depending on the specific task at hand. However, we introduce a “high performance” measure to estimate roughly if the strategies positively or negatively affect the performances. For instance, model compression techniques often lower model performance, whereas incremental and meta-learning typically boost it. As a result, we use six distinct measures to evaluate the techniques: computational efficiency, memory footprint, low energy consumption, quick training time, reduced bandwidth, and high performance. Table II compares the various families of techniques against these requirements.
The outcomes, as depicted in Table II, should be interpreted as informal and approximate assessments aimed at providing a broad understanding of the general strengths and weaknesses of the compared techniques within the context of edge learning. A checkmark (”✓”) implies that, in general, the technique offers assistance or advantages when applied to edge learning with respect to the specified requirement. Conversely, a cross mark (”✗”) indicates that, in general, the requirement represents a weakness of the technique in the edge learning context. It is important to note that techniques like FL, quantization, and others have various variants and specific approaches that influence how these methods align with the requirements. For instance, while quantization techniques may result in a minor decrease in performance in most scenarios [226], leading to an ”✗” check in the ”High Performance” column. Certain specific approaches to quantization may exceptionally yield no performance loss or even improvements, as demonstrated in studies such as [190]. Therefore, it is important to note that the assessment provided by these symbols should be considered as rough estimates, as the effectiveness of a technique, and how it fares against a specific requirement, can vary depending on diverse factors such as variant versions, implementation details, use cases, tasks, and hardware platforms. Accordingly, table II offers a general overview rather than a definitive judgment on the suitability of each technique for every situation.